A sound wave of the form s = sm cos(kx - ωt + φ) travels at 343 m/s through air in a long horizontal tube. At one instant, air molecule A at x = 2.04 m is at its maximum positive displacement of 6.20 nm and air molecule B at x = 2.08 m is at a positive displacement of 2.00 nm. All the molecules between A and B are at intermediate displacements. What is the frequency of the wave?
Answer #1

Maximum positive displacement, Xm = 2/6.2 = 0.323

K[Xb - Xa] = arccos(0.323)

K[2.08 - 2.04] = 1.24 rad

(2 pi/lambda) (0.04) = 1.24 rad

Lambda = 0.203 m

Frequency, f = 343/0.203 = 1692.29 Hz


answered by: ANURANJAN SARSAM